OpenStudy (anonymous):

A line passes through the point −2, 6 and has a slope of −5/2 . Write an equation in slope-intercept form for this line.

OpenStudy (anonymous):

@optiquest

OpenStudy (anonymous):

Substitute your values into this equation{y=mx+b} put your slope where the m is {-5/2] as m= slope and your y value into the b slot {6} and you should get y= -5/2+6{hope this helps:).)